A piston-cylinder device initially contain 0.4 m3 of 1.8-kg saturated liquid water and steam mixture at 110 kPa. A electric heater in the device is turned on and passes a current of 1.7 A from a 120-V source. Only transfer of electrical work and boundary work exist. The piston initially rests on a set of stops, and a pressure of 200 kPa is required to move the piston. Afterwards the system continues to raise the piston and eventually reaches the temperature of 400 Celcius.

Find the initial quality of the system,initial internal energy (kj/kg), time it takes for the system to just move the piston (hours), final volume of system (m3), and total time it takes for the system to reach its final state (hours)
Answer #1

200 kPa V.0.4m² mas 1.8 kg 6 w = (1.7x120) = 204 walt initial Conditions: saturated liquid water & steam mixture. volume v=0.G initial quality of steam V (of + at 110 kPa. 0-2224 0.001046 + * [1.5492 -0.001046] x = 0.1428 2 initial internal Energy Uat 200 kpa, (2 bar) [from steam tables) VF 0.001061 m²/kg m/g va = 0.88540 ht 504.7 kJ /kg hfg 2201.6 kJ/kg (a + 22 Mfgo) e 2:heat required to just raise the Piston in (₂-0) 1.8 (10112 253 - 726.154) W = 513.178 KJE. Power given by electric heater 1.after that, System Continues to raise the Piston and eventually reaches the temperature 400°c. 2-3 is Constant Pressure Proce6. Total time it takes for System to reach its final state T= time taken for 1-2 + time taken for 2-3 time taken for 2-3 ProcI hope you will understand the answer and i gave it in a detailed manner and pls download steam tables to solve this problem for reference and please thumbs up...